# SysCaps: Language Interfaces for Simulation Surrogates of Complex Systems

### v2

@@ 1 Introduction
Data-driven surrogates enable computational scientists to efficiently predict the results of expensive numerical simulations that run on supercomputers (Lavin et al., 2021; Carter et al., 2023). Surrogates are particularly valuable for emulating simulations of complex energy systems (CES), which model dynamic interactions between humans, earth systems, and infrastructure. Examples of CES include buildings (Vazquez-Canteli et al., 2019; Dai et al., 2023; Bhavsar et al., 2023), electric vehicle fleets (Vepsäläinen et al., 2019), and microgrids (Du & Li, 2019). Advancing the science of CES contributes to reducing emissions and accelerating the adoption of clean energy, which is needed to address the impacts of climate change.
These models perform a fairly standard regression task, predicting simulation output quantities of interest from a) an input system configuration and b) a deployment scenario. For example, we might want to predict the amount of energy a building will consume given a) a list of building characteristics and b) a weather timeseries spanning an entire year. In this case, this involves performing long sequence timeseries regression, which traditional regression techniques such as gradient-boosted decision trees have difficulty with (Bhavsar et al., 2023; Zhang et al., 2021).
@@
In general, it is unknown whether textual inputs, and particularly natural language inputs, are suitable for real-world tabular regression tasks. Tabular data, such as the system attributes in question, are sets of both discrete (categorical, binary, or string) and continuous (numeric) variables. Previous work demonstrated inconclusive evidence when using pretrained language models to do tabular regression from text-encoded inputs, with and without modifications to the architecture (Dinh et al., 2022; Jablonka et al., 2024; Bellamy et al., 2023; Yan et al., 2024), motivating further study. Regression with text-encoded tabular inputs is promising because a) language is a more intuitive and flexible user interface than traditional encoding strategies (e.g., one-hot encodings), and b) using pretrained language embeddings to encode system attributes unlocks the ability to exploit the semantic information contained in SysCaps to generalize across related systems.
@@
  • We introduce a simple and lightweight multimodal surrogate model architecture for timeseries regression that a) fuses text embeddings obtained from pretrained language models (LMs) with b) timeseries encoded by a bidirectional sequence encoder. We expect this to be insightful for future multimodal text and timeseries studies.
@@
Our experiments are based on two real-world CES simulators of buildings and wind farm wake. We rigorously evaluate accuracy on held-out systems and show that SysCaps-augmented surrogates have better accuracy than one-hot baselines. We also show generalization beyond the capabilities of traditional regression approaches enabled by the use of text embeddings, e.g., robustness to replacing attributes names with synonyms in test captions. We qualitatively show that text interfaces unlock system design space exploration via language. As there are no standard benchmarks for comparing surrogate modeling performance for CES, we will open-source all data and code and contribute the generated SysCaps datasets to facilitate future work.
@@ 2 Related Work
Multimodal text and timeseries forecasting: Timeseries forecasting aims to predict future values of an input timeseries given past values. Surrogate modeling can be cast as a forecasting problem when the goal is to train a model to emulate a dynamical system and predict its future behavior (e.g., predicting future energy demand from past energy usage). Previous work has explored multimodal timeseries forecasting where auxiliary text data is introduced as covariates to improve the forecasting accuracy (Rodrigues et al., 2019; Emami et al., 2023a; Jin et al., 2024). Notably, Time-LLM (Jin et al., 2024) “reprograms” an LLM to process both text prompts and timeseries. However, our timeseries regression setting (Section 3) differs in that our surrogates are trained to map simulator inputs (e.g., building characteristics and a weather timeseries) to simulator outputs (e.g., energy usage). In our problem, models critically depend on the system information encoded as text, whereas in Time-LLM the text only contains auxiliary information that slightly improves forecasting accuracy. This critical dependence partially motivates our development of a custom architecture for fusing text and timeseries embeddings. Recent evidence (Merrill et al., 2024; Tan et al., 2024) also brings into question whether LLMs are useful at all for reasoning about temporal data; thus, research on lightweight multimodal models is warranted.
@@
Knowledge-enhanced PDE surrogates: Numerical simulation of partial differential equations (PDEs) is extremely computationally intensive, and thus a large body of SciML work is focused on developing neural PDE surrogates which are efficient to evaluate. Recent work has tried to encode knowledge about the physical system under study (the PDE) into a surrogate to facilitate generalization within and across families of PDEs. Specifically, methods that embed equation parameters (i.e., the system attributes) within the architecture to generalize to unseen parameters include CAPE (Takamoto et al., 2023) and those explored in Gupta & Brandstetter (2022). Others embed structural knowledge about the PDE into the surrogate model architecture (Rackauckas et al., 2020; Ye et al., 2024) or the loss (Raissi et al., 2019). Concurrent work has explored “PDE captions” (Lorsung et al., 2024; Yang et al., 2024), which are a type of SysCaps for neural PDE surrogates where the system knowledge is PDE equations encoded as text.
@@ 3 Problem Statement
To summarize, we have a timeseries regression problem modeled as $y_{1:T}=f(x_{1:T},z)$. By conditioning the surrogate on system knowledge $z$, it can potentially generalize to new system configurations. However, learning transferable representations of variable-length, heterogeneous input features such as $z$ is notoriously difficult for deep neural networks, and is a key focus of tabular deep learning (see survey  (Badaro et al., 2023)). In our work, we develop and analyze a framework for learning multimodal surrogates where $z$ is encoded as text.
@@ 5 Text and Timeseries Surrogate Model
Although more sophisticated loss functions than Eq. 1 could be used that account for predictive uncertainty, we left this extension for future work to simplify our exposition and experiments.
Multimodal encoder $g_{\psi}$: A text encoder $g_{\psi}^{\mathsf{text}}$ extracts an embedding $\hat{z}$ from a SysCap $z$ then broadcasts and concatenates this embedding with the timeseries inputs to create a sequence of multimodal feature vectors. These features get processed by a bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$ to produce a sequence of time-dependent fused multimodal features $e_{1:T}$, $e_{1:T}=g_{\psi}^{\mathsf{seq}}(g_{\psi}^{\mathsf{text}}(z),x_{1:T})$ , which are finally used to regress outputs.
@@ 6 Experiments
This section presents our experimental results, which include: quantifying the quality of LLM-generated SysCaps (Section 6.1), accuracy on held-out systems (Section 6.2), generalization under distribution shifts (Section 6.3), a design space exploration application (Section 6.4), and a study on SysCaps prompt augmentation (Section 6.5). These experiments are based on two real-world CES simulators for buildings and wind farms. All SysCaps are synthetically generated in this work. We provide additional qualitative examples of SysCaps in Appendix B.
Setup: Our main experiments focus on training building stock surrogate models for the building energy simulator EnergyPlus (Crawley et al., 2001). Given an annual hourly weather timeseries ($T$ = 8,760) with 7 variables and a list of tabular building attributes, surrogates predict the building’s energy consumption at each hour of the year. Each building initially has 17 attributes; using RFE with a tuned LightGBM (Ke et al., 2017) model, we selected the 13 most important attributes. We use the commercial building split of the Buildings-900K dataset (Emami et al., 2023b), which are building stock simulation runs representative of the United States commercial building stock. Commercial building stock surrogates can provide significant speedups compared to EnergyPlus, e.g., 96% (Zhang et al., 2021). Since this dataset only provides energy timeseries, we extract the building configuration and weather timeseries from the End-Use Load Profiles database (Wilson et al., ) for each building. Our training set is comprised of 330K buildings, and we use 100 buildings for validation, and 6K held-out buildings for testing. We also reserved a held-out set of 10K buildings for RFE. We carefully tune the hyperparameters of all models (details in Appendix A.2).
@@ 6.2 Accuracy On Held-Out Systems
Following Emami et al. (2023b), we use the normalized root mean square error (NRMSE) metric to compare model accuracy, averaged across 3 random seeds. In addition to comparing the ResNet MLP, LSTM, and SSM, we also trained a tuned LightGBM baseline. Does the sequential architecture matter? Yes—Table 1 shows that the LSTM and SSM encoders outperform both the ResNet and our carefully tuned LightGBM baseline, and the SSM outperforms the LSTM. How do different system attribute encoding approaches compare? Surprisingly, the best SSM and LSTM with key-value SysCaps achieve comparable building-hourly test accuracy to one-hot baselines and the best accuracy at the stock-annual aggregation level (equivalent to the normalized mean bias error—see Appendix A.1). The SSM and LSTM with natural language SysCaps has slightly worse accuracy than with the key-value SysCaps, yet they comfortably outperform the non-sequential models (including LightGBM) and the one-hot baselines at the stock-annual aggregation level. We initially expected to see a non-negligible drop in regression accuracy for SysCaps models, even key-value SysCaps, because the text encoder has to compress the caption into a single embedding vector. However, the BERT encoder is sufficiently expressive to mitigate this. We observe that the stock-level NRMSE reduces by about half when switching from DistilBERT to BERT. We believe the performance gap between key-value templates and natural language SysCaps is mostly explained by the caption quality (Table 3). We ablate the importance of the system attributes by training an SSM baseline with these inputs removed (SSM with Attribute Encoding “X”); this model is unable to learn this task.
@@ 6.3 Caption Generalization
Length generalization: We assess how accuracy varies when surrogates are provided with natural language SysCaps having different lengths than seen during training. We evaluate zero-shot generalization to the short and long captions. The results (Table 3) show a small increase in error for shorter captions with a larger increase in error for longer captions, as might be expected. The error on long captions remains lower than the error achieved by our tuned LightGBM baseline.
@@ 6.5 Prompt Augmentation: Wind Farm Wake
Table 5: Wind farm surrogate accuracy. The base architecture is ResNet. Average across 3 random seeds.
@@ 7 Conclusion
In this work, we introduced a lightweight, multimodal text and timeseries surrogate models for complex energy systems such as buildings and wind farms, and described a process for using LLMs to synthesize natural language descriptions of such systems, which we call SysCaps. Our experiments showcase SysCaps-augmented surrogates that achieve better accuracy than standard feature engineering (e.g., one-hot encoding) while also enjoying the advantages of using text embeddings such as robustness to caption paraphrasing (e.g., synonyms of attributes). For a problem with only a small number of training systems available, we showed that SysCaps prompt augmentation has a regularizing effect that helps mitigate overfitting. Overall, these results underscore that language is a viable interface for interacting with real-world surrogate models.
@@
Broader impacts: Improving surrogate models for CES has the potential to accelerate the transition to cleaner energy sources. To avoid unfair outcomes from decisions made with CES surrogates, care should be taken when deciding what simulation runs to use as training data and when selecting an LLM to use for caption generation.
Future work: A future extension of this work might explore how to use language to also interface with the timeseries simulator inputs, possibly through summary statistics. For example, to study how the complex system behaves when the average exogenous temperature is increased by five degrees. Another promising extension is to conduct a series of human studies with domain scientists to evaluate the quality of SysCaps in the context of, e.g., system design optimization. Moreover, an important question is how we might create surrogate foundation models that generalize not only across system configurations for a single simulator, but also generalize across different simulators.
@@ A.1 Metrics
The stock-annual NRMSE, is where the NRMSE is normalized by the annual stock energy consumption:
@@
We use the AdamW Loshchilov & Hutter (2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$ = 1e-9, and weight decay of 0.01 for all experiments. The early stopping patience is 50 for all experiments. All models are trained with a single NVIDIA A100-40GB GPU. The longest training runs take 1-2 days and the shortest 2-3 hours.
@@ A.2 Hyperparameters
Table 6: Buildings: hyperparameters.
@@
Table 7: Wind: hyperparameters.
@@
See Table 6 and Table 7.
@@ A.2.2 Wind farm
Appendix B Additional SysCaps examples
In this section we visualize examples of prompts and the corresponding SysCaps outputs for the buildings simulator (Figures 6, 7, and 8) and the wind farm simulator (Figures 9, 10, and 11).

### v3

@@ 1 Introduction
Data-driven surrogates enable computational scientists to efficiently predict the results of expensive numerical simulations that run on supercomputers (Lavin et al., 2021; Carter et al., 2023). Surrogates are particularly valuable for emulating simulations of complex energy systems (CES), which model dynamic interactions between humans, earth systems, and infrastructure. Examples of CES include buildings (Vazquez-Canteli et al., 2019; Dai et al., 2023; Bhavsar et al., 2023), electric vehicle fleets (Vepsäläinen et al., 2019), and microgrids (Du & Li, 2019). Advancing the science of CES contributes to efforts aimed at reducing emissions and improving the resiliency of power systems.
These surrogates perform a fairly standard regression task, predicting simulation output quantities of interest from a) an input system configuration and b) a deployment scenario. For example, we might want to predict the amount of energy a building will consume given a) a list of building characteristics and b) a weather timeseries spanning an entire year. In this case, this involves performing long sequence timeseries regression, which traditional regression techniques such as gradient-boosted decision trees have difficulty with (Bhavsar et al., 2023; Zhang et al., 2021).
@@
In general, it is unknown whether textual inputs, and particularly natural language inputs, are suitable for real-world tabular regression tasks. Tabular data, such as the system attributes in question, are sets of both discrete (categorical, binary, or string) and continuous (numeric) variables. Previous work demonstrated inconclusive evidence when using language models to do tabular regression from text-encoded inputs, with and without modifications to the architecture (Dinh et al., 2022; Jablonka et al., 2024; Bellamy et al., 2023; Yan et al., 2024), motivating further study. Regression with text-encoded tabular inputs is promising because a) language is a more intuitive and flexible user interface than traditional encoding strategies (e.g., one-hot encodings), and b) using language embeddings to encode system attributes unlocks the ability to exploit the semantic information contained in SysCaps to generalize across related systems.
@@
  • We introduce a simple and lightweight multimodal surrogate model architecture for timeseries regression that a) fuses text embeddings obtained from fine-tuned language models (LMs) with b) timeseries encoded by a bidirectional sequence encoder. We expect this to be insightful for future multimodal text and timeseries studies.
@@
Our experiments are based on two real-world CES simulators of buildings and wind farm wake. We rigorously evaluate accuracy on held-out systems and show that SysCaps-augmented surrogates have better accuracy than one-hot baselines. We also show generalization beyond the capabilities of traditional regression approaches enabled by the use of text embeddings, e.g., robustness to replacing attributes names with synonyms in test captions. We qualitatively show that text interfaces unlock system design space exploration via language. As there are no standard benchmarks for comparing surrogate modeling performance for CES, we open-source all code and data at https://github.com/NREL/SysCaps to facilitate future work.
@@ 2 Related Work
Multimodal text and timeseries forecasting: Timeseries forecasting aims to predict future values of an input timeseries given past values. Surrogate modeling can be cast as a forecasting problem when the goal is to train a model to emulate a dynamical system and predict its future behavior (e.g., predicting future energy demand from past energy usage). Previous work has explored multimodal timeseries forecasting where auxiliary text data is introduced as covariates to improve the forecasting accuracy (Rodrigues et al., 2019; Emami et al., 2023a; Jin et al., 2024). Notably, Time-LLM (Jin et al., 2024) “reprograms” an LLM to process both text prompts and timeseries. However, our timeseries regression setting (Section 3) differs in that our surrogates are trained to map simulator inputs (e.g., building characteristics and a weather timeseries) to simulator outputs (e.g., energy usage). In our problem, models critically depend on the system information encoded as text, whereas in Time-LLM the text only contains auxiliary information that slightly improves forecasting accuracy. This critical dependence partially motivates our development of a lightweight architecture for fusing text and timeseries embeddings. Also, recent evidence (Merrill et al., 2024; Tan et al., 2024) brings into question whether LLMs are useful at all for reasoning about temporal data.
@@
Knowledge-enhanced PDE surrogates: Numerical simulations of partial differential equations (PDEs) are computationally intensive; thus, much SciML aims to train PDE surrogates which are efficient to evaluate on unseen inputs. Recent work tries to encode knowledge about the PDE into the surrogate to facilitate generalization within and across families of PDEs. Specifically, CAPE (Takamoto et al., 2023) and methods explored in Gupta & Brandstetter (2022) embed equation parameters (i.e., the system attributes) within the architecture to generalize to unseen parameters. Others embed structural knowledge about the PDE into the surrogate model architecture (Rackauckas et al., 2020; Ye et al., 2024) or the loss (Raissi et al., 2019). Concurrent work has explored “PDE captions” (Lorsung et al., 2024; Yang et al., 2024), which are a type of SysCaps for neural PDE surrogates where the system knowledge is PDE equations encoded as text.
@@ 3 Problem Statement
To summarize, we have a timeseries regression problem modeled as $y_{1:T}=f(x_{1:T},z)$. By conditioning the surrogate on system knowledge $z$, it can potentially generalize to new system configurations. However, learning transferable representations of variable-length, heterogeneous input features such as $z$ is notoriously difficult for deep neural networks, and is a key focus of tabular deep learning (see survey by Badaro et al. (2023)). In our work, we develop and analyze a framework for learning multimodal surrogates where $z$ is encoded as text.
@@ 5 Text and Timeseries Surrogate Model
Although more sophisticated loss functions than Eq. 1 could be used to account for predictive uncertainty, we left this extension for future work to simplify our exposition and experiments.
Multimodal encoder $g_{\psi}$: A text encoder $g_{\psi}^{\mathsf{text}}$ extracts an embedding $\hat{z}$ from a SysCap $z$, then broadcasts and concatenates this embedding with the timeseries inputs to create a sequence of multimodal feature vectors. These features get processed by a bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$ to produce a sequence of time-dependent fused multimodal features $e_{1:T}$, $e_{1:T}=g_{\psi}^{\mathsf{seq}}(g_{\psi}^{\mathsf{text}}(z),x_{1:T})$ , which are finally used to regress outputs.
@@ 6 Experiments
This section presents our experimental results on two real-world CES simulators for buildings (Section 6.1-6.4) and wind farms (Section 6.5). Our experiments study the quality of LLM-generated SysCaps (Section 6.1), accuracy on held-out systems (Section 6.2), generalization under distribution shifts (Section 6.3), show a design space exploration application (Section 6.4), and examine SysCaps prompt augmentation (Section 6.5). All SysCaps are synthetically generated in this work. We provide additional qualitative examples of SysCaps in Appendix B.
Building stock simulation data: For the main experiments in Section 6.1-6.4 we train building stock surrogate models for the building energy simulator EnergyPlus (Crawley et al., 2001). Given an annual hourly weather timeseries ($T$ = 8,760) with 7 variables and a list of tabular building attributes, surrogates predict the building’s energy consumption at each hour of the year. Each building initially has 17 attributes which we reduce to 13 with RFE and LightGBM (Ke et al., 2017). We use commercial buildings from the Buildings-900K dataset (Emami et al., 2023b).Commercial building stock surrogates can provide significant speedups compared to EnergyPlus, e.g., 96% (Zhang et al., 2021). Since this dataset only provides energy timeseries, we manually extracted the building configuration and weather timeseries from the End-Use Load Profiles database (Wilson et al., ) for each building. Our training set is comprised of 330K buildings, and we use 100 buildings for validation and 6K held-out buildings for testing. We also reserved a held-out set of 10K buildings for RFE. We carefully tune the hyperparameters of all models (details in Appendix A.2).
@@ 6.2 Accuracy On Held-Out Systems
In this section, we compare the accuracy of SysCaps surrogate models that vary by textual attribute encoding (key-value (-kv) and natural language (-nl)), text encoder (BERT, DistilBERT), and sequence encoder (LSTM, SSM). We also train SysCaps surrogates with a non-sequential encoder (ResNet), without attribute inputs, and without text encoder fine-tuning, as ablations. Baselines are a tuned LightGBM Gradient Boosting Decision Tree and a ResNet, LSTM, and SSM surrogate with onehot encoded attributes. Following Emami et al. (2023b), we use the normalized root mean square error (NRMSE) metric averaged across 3 random seeds.
Does the sequential architecture matter? Yes—Table 1 shows that the LSTM and SSM encoders outperform both the ResNet and our carefully tuned LightGBM baseline, and the SSM outperforms the LSTM. How do different system attribute encoding approaches compare? Surprisingly, SysCaps-kv achieves the best accuracy overall at the stock-annual aggregation level (equivalent to the normalized mean bias error—see Appendix A.1) and comparable building-hourly accuracy to one-hot baselines. The SysCaps-nl models have slightly worse accuracy than SysCaps-kv, yet they comfortably outperform the non-sequential models (including LightGBM) and the one-hot baselines at the stock-annual aggregation level. We initially expected to see a non-negligible drop in regression accuracy for SysCaps models, even key-value SysCaps, because the text encoder has to compress the caption into a single embedding vector, which causes information loss. However, the BERT encoder is expressive enough to mitigate this. We observe that stock-annual NRMSE reduces by about half when switching from DistilBERT to BERT. We believe the gap between key-value and natural language is mostly explained by caption quality (Table 3). We ablate the importance of using attributes by training an SSM baseline without attributes (Attribute Encoding “X”). This model is unable to learn this task. We trained a SysCaps-nl model without fine-tuning BERT, but it does poorly (stock-annual NRMSE of 0.356), showing the importance of fine-tuning for multi-modal alignment.
@@ 6.3 Caption Generalization
Length generalization: We assess how accuracy varies when surrogates are provided with natural language SysCaps having different lengths than seen during training. We evaluate zero-shot generalization to the short and long captions. The results (Table 3) show a small increase in error for shorter captions with a larger increase in error for longer captions. However, the error on long captions remains lower than the error achieved by our tuned LightGBM baseline.
@@ 6.5 Prompt Augmentation: Wind Farm Wake
Table 5: Wind farm surrogate accuracy. The base architecture is ResNet. Average across 3 random seeds. SysCaps-nl* does not use prompt augmentation.
@@ 7 Conclusion
In this work, we introduced a lightweight, multimodal text and timeseries surrogate models for complex energy systems such as buildings and wind farms, and described a process for using LLMs to synthesize natural language descriptions of such systems, which we call SysCaps. Our experiments showcase SysCaps-augmented surrogates that achieve better accuracy than standard feature engineering (e.g., one-hot encoding) while also enjoying the advantages of using text embeddings such as robustness to caption paraphrasing (e.g., synonyms of attributes). For a problem with only a small number of training systems available, we showed that SysCaps-nl prompt augmentation has a regularizing effect that helps mitigate overfitting. Overall, these results underscore that language is a viable interface for interacting with real-world surrogate models.
@@
Future work: A future extension of this work might explore how to use language to also interface with the timeseries simulator inputs, possibly through summary statistics. For example, to study how the complex system behaves when the average exogenous temperature is increased by five degrees. It is natural to expect that non-experts may benefit more from our approach if the LLM is also instructed to simplify the simulator metadata or to provide explanations of technical concepts. Conducting interactive evaluations with non-experts will be important to obtain feedback for further improving the approach. Likewise, conducting a series of studies with domain scientists to evaluate the quality of SysCaps in the context of, e.g., system design optimization is promising. We did not conduct user studies in this work, as we first aimed to establish technical feasibility of this surrogate modeling approach. Finally, an important question is how we might create surrogate foundation models that generalize not only across system configurations for a single simulator, but also generalize across different simulators.
@@ A.1 Metrics
The stock-annual NRMSE is where the NRMSE is normalized by the annual stock energy consumption:
@@
We use the AdamW (Loshchilov & Hutter, 2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, $\epsilon$ = 1e-9, and weight decay of 0.01 for all experiments. The early stopping patience is 50 for all experiments. All models are trained with a single NVIDIA A100-40GB GPU. The longest training runs take 1-2 days and the shortest 2-3 hours.
@@ A.2 Hyperparameters
Table 6: Building stock surrogate model hyperparameters.
@@
Table 7: Wind farm wake surrogate model hyperparameters.
@@
See Table 6 for hyperparameter sweep details for the buildings experiments and Table 7 for hyperparameter sweep details for the wind farm experiments.
@@ A.2.1 Buildings
Multi-class attribute classifier: We implement the classifier on top of the text encoder by adding a linear layer for each attribute type, where this layer predicts logits for each attribute’s classes. We use AdamW with a learning rate of 3e-4, early stopping with patience 5, batch size 128, and max epochs 100. We do not freeze the text encoder weights.
@@ A.2.2 Wind farm
Appendix B SysCaps Prompts
In this section we visualize examples of prompts and the corresponding SysCaps LLM outputs for the buildings simulator (Figures 6, 7, and 8) and the wind farm simulator (Figures 9, 10, and 11).
@@ Appendix B SysCaps Prompts
Appendix C Qualitative Examples of Test Predictions
We visualize in Figure 12 the predicted energy timeseries for one test building by the SysCaps-nl and SysCaps-kv models alongside the weather timeseries and each SysCaps type (key-value, short, medium, and long). We also show test predictions for a wind farm (Figure 13) alongside the key-value SysCaps and each style-augmented natural language SysCaps.
Figure 12: (Top) Predicted hourly energy consumption for a randomly selected week. (Second from top) We visualize 1 out of 7 input weather timeseries, Dry Bulb Temperature. The text boxes show the medium, short, long, and key-value SysCaps created for this particular test building.
Figure 13: We show the LLM-generated SysCaps using four different styles as a prompt augmentation strategy for test wind farm layout # 214. The key-value SysCaps is shown at the bottom. Model predictions are shown at the top next to the true ground truth value.